a man pushes a 2,220 kg car with an average force of 128 N for 7.56 s. The impulse on the car is ___ N•s
1
Expert's answer
2019-02-06T12:51:47-0500

By the definition of the impulse, we have:

"J = F \cdot \Delta t,"

here,

"J"

is the impulse on the car,

"F"

is the average force that acts on the car,

"\Delta t"

is the time during which the force acts on the car.

Then, we get:

"J = F \cdot \Delta t = 128 N \cdot 7.56 s = 968 N \cdot s."

Answer:

"J = 968 N \cdot s"

.